### v1

# History, Development, and Principles of Large Language Models—An Introductory Survey

Abstract
Language models serve as a cornerstone in natural language processing (NLP), utilizing mathematical methods to generalize language laws and knowledge for prediction and generation. Over extensive research spanning decades, language modeling has progressed from initial statistical language models (SLMs) to the contemporary landscape of large language models (LLMs). Notably, the swift evolution of LLMs has reached the ability to process, understand, and generate human-level text. Nevertheless, despite the significant advantages that LLMs offer in improving both work and personal lives, the limited understanding among general practitioners about the background and principles of these models hampers their full potential. Notably, most LLMs reviews focus on specific aspects and utilize specialized language, posing a challenge for practitioners lacking relevant background knowledge. In light of this, this survey aims to present a comprehensible overview of LLMs to assist a broader audience. It strives to facilitate a comprehensive understanding by exploring the historical background of language models and tracing their evolution over time. The survey further investigates the factors influencing the development of LLMs, emphasizing key contributions. Additionally, it concentrates on elucidating the underlying principles of LLMs, equipping audiences with essential theoretical knowledge. The survey also highlights the limitations of existing work and points out promising future directions.
1. INTRODUCTION
Language stands as humanity’s most potent tool, enabling the expression of thoughts and emotions and facilitating communication with others (Pinker, 2023; Turing, 2009). However, machines lack the intrinsic ability to comprehend and communicate in human language, necessitating powerful Artificial Intelligence (AI) algorithms to acquire such capabilities. The ultimate aim of AI research is the creation of machines capable of reading, writing, and conversing like humans (Dwivedi et al., 2023; Zhao et al., 2023; Jin et al., 2023). Natural language processing (NLP) emerges as the AI branch dedicated to realizing this goal. At the core of NLP are language models, tasked with predicting or generating the probability or likelihood of linguistic units (e.g., words, phrases, sentences) based on context. The evolutionary stages of language models unfold chronologically from the initial statistical language models (SLMs) to subsequent neural language models (NLMs), progressing to pre-trained language models (PLMs), and reaching the current state of large language models (LLMs). Specifically, SLMs employ simple probability distributions to model word sequences, while NLMs utilize neural networks to grasp complex patterns and representations of language. PLMs leverage large-scale corpora and self-supervised learning to capture general linguistic knowledge, and LLMs extend PLMs by incorporating massive data, computation, and algorithms, resulting in more expressive, sophisticated, and adaptable language models (Fu et al., 2022;Wei et al., 2022).
In recent years, significant advancements in NLP have been achieved with LLMs, such as the generative pre-trained transformer (GPT) series of models (Radford et al., 2018; Radford et al., 2019; Brown et al., 2020; OpenAI, 2023). Extensively trained on textual data, these models demonstrate the ability to generate human-level texts and perform language-based tasks with exceptional precision. Despite these advancements having greatly benefited various aspects of work and life, LLMs might not be widely understood by practitioners, particularly those without a background in NLP. To provide practitioners with a fundamental understanding of LLMs, this review introduces it across six dimensions: history, development, principles, applications, drawbacks and future directions. This survey distinguishes itself from other reviews on PLMs or LLMs (Zhao et al., 2023; Liu et al., 2023; Han et al., 2021; Shanahan, 2022), which often concentrate on specific aspects using technical language; this survey aims to elucidate LLMs and their principles in a more accessible manner, intending to maximize its full potential.
The subsequent sections of this review are structured as follows: Section 2 details the history of LLMs and the contributing factors behind their rapid growth, covering the evolution of language models, increased data diversity, computational advancements, and algorithmic innovations. Section 3 offers an overview of LLMs principles, using the GPT family of models as an accessible example for better comprehension. Moving on, Section 4 delves into the diverse applications of LLMs, spanning drug discovery, finance, medical, and legal domains. Section 5 critically examines the drawbacks of current state-of-the-art LLMs, and finally, Section 6 concludes the survey.
2. HISTORY AND DEVELOPMENT OF LARGE LANGUAGE MODELS
In this section, we will explore the history of the language model and analyze the developmental dynamics and influences shaping LLMs.
2.1 History of LLMs
To better understand LLMs, this section will follow the developmental stages of language models and introduce SLMs, NLMs, PLMs, and LLMs. Figure 1 provides a visual map of the history of the development of the language model.
Figure 1: History and development of language models.
  • Statistical Language Model: SLMs (Jelinek, 1998; Stolcke, 2002; Rosenfeld, 2000) originated in the 1990s as mathematical models addressing contextually relevant properties of natural language from a probabilistic statistical perspective. The essence of statistical language modeling lies in ascertaining the probability of a sentence occurring within a text. Considering $S$ as the sentence “I am very happy”, $P(\omega_{i})$ signifies the probability of the $i$-th word in the sentence: $\omega_{1}$ as “I”, $\omega_{2}$ as “am”, $\omega_{3}$ as “very”, and $\omega_{4}$ as “happy”. Now, the objective is to ascertain the likelihood of $S$ appearing in the text, denoted as $P(S)$:
    $P(S)=P(\omega_{1},\omega_{2},\omega_{3},\omega_{4})=P(I,am,very,happy)$ (1)
    To calculate this probability, the conditional probability can be employed:
    $P(I,am,very,happy)=P(I)\cdot P(am\mid I)\cdot P(very\mid I,am)\cdot P(happy\mid I,am,very)$ (2)
    where $P(I)$ represents the probability of the word “I” appearing and $P(am|I)$ stands for the probability of “am” appearing given that “I” has appeared. When we multiply $P(am|I)$ by $P(I)$, it fulfills the condition of “I” appearing in $P(am|I)$, resulting in the probability of “I am” appearing together as $P(I,am)=P(I)\cdot P(am|I)$. Now, the question arises: how do we calculate the conditional probability of the occurrence of each word? The answer lies in Maximum Likelihood Estimation, enabling us to estimate probabilities by substituting them with frequencies when the sample size is sufficiently large, given by:
    $P(w_{i}|w_{1}w_{2}\cdots w_{i-1})=\frac{P(w_{1}\cdots w_{i-1}w_{i})}{P(w_{1}w_{2}\cdots w_{i-1})}=\frac{{\bf C}(w_{1}w_{2}\cdots w_{i})}{{\bf C}(w_{1}w_{2}\cdots w_{i-1})}$ (3)
    where $\mathbf{C(\cdot)}$ represents the count of occurrences of the subsequence in the training set. Using this formula, we can calculate the likelihood of each word as the $i$-th word given the preceding $i-1$ words. Then, we select the $i$-th word by choosing the word associated with the highest probability. The previous discussion assumes that the $n$-th word is related to the initial $n-1$ words, consistent with our intuition and commonly referred to as the n-gram model. In order to efficiently compute conditional probabilities, it is necessary to pre-compute and save $C(X)$required for the conditional probability computation, where $X$ is a sentence of length $n$. The number of possible sentences $X$ grows exponentially with the size of the vocabulary. For instance, with $1000$ different words, there exist ${1000}^{n}$ potential sequences of length $n$. However, excessively large values of $n$ pose storage limitations. Typically, $n$ is confined to $2$ or $3$, causing each word to relate to only its first $1$ or $2$ preceding words, ultimately leading to a reduction in the model’s accuracy.
  • Neural Language Models: NLMs (Bengio et al., 2000; Mikolov et al., 2010; Kombrink et al., 2011) leverage neural networks to predict the probabilities of subsequent words within sequences. They effectively handle longer sequences and mitigate the limitations associated with small $n$ in SLMs. Before delving into neural networks, let’s grasp the concept of word vectors. Humans effortlessly comprehend word meanings. For instance, “cat” and “dog” share closer semantic connections than “cat” and “tree” since they both represent animals. But how does a computer accomplish this? Computers operate using binary code — 0s and 1s — so human language needs translation into binary. Word vectors can accomplish this, which are numerical representations of human language, where each word corresponds to a distinct vector. These vectors usually possess fixed dimensions and can simulate word relationships through the angles between them. An intriguing example is the angle between the word vectors for “cat” and “dog” which is smaller than the angle between the word vector for “cat” and “tree”. Another illustrative example is “$M\ essi-Argentina+P\ ortugal=Cristiano\ Ronaldo$”. Word2Vec (Mikolov et al., 2013) is a widely recognized tool for computing word vectors. Its function involves converting words into dense numerical representations, making words with similar semantics closer together in the vector space.
    Figure 2: Neural language model.
    The efficacy of Word2Vec is notably influenced by its utilization of neural networks, which mirrors the structure and functionality of the human brain, comprising interconnected simple units known as neurons organized into different layers. Figure 2 provides a visualization of a simple NLM structure, composed of input, hidden, and output layers. Within this structure, each word vector $x_{t}\in\mathbb{R}^{m}$, the matrix $W_{xh}\in\mathbb{R}^{m(n-1)\times h}$, and the matrix $W_{ho}\in\mathbb{R}^{h\times|V|}$, where $h$ signifies the number of neurons in the hidden layer, and $V$ is the vocabulary containing all words recognized by the model. NLMs operate akin to the n-gram concept, assuming that the probability of each word depends only on its previous n-1 words. The first layer—the input layer—concatenates the word vectors of n-1 words, forming a vector $X\in\mathbb{R}^{m\times(n-1)}$. Subsequently, the second layer—the hidden layer—derives the output $H$ by applying a non-linear activation function, like sigmoid or tanh, to the matrix product of $X$ and $W_{xh}$. Following this, the third layer—the output layer—aims to forecast the subsequent word based on the hidden layer’s output. With $|V|$ neurons within the output layer, the resulting output vector $O\in\mathbb{R}^{|V|}$ is computed by multiplying $H$ and $W_{ho}$. This resultant vector then undergoes the Softmax function, producing a vector $O’$ ’ containing the probability value assigned to each word. The Softmax function, concerning the output of the $i$-th neuron, is defined as follows:
    $O_{i}^{\prime}=\operatorname{Softmax}\left(O’_{i}\right)=\frac{\exp\left(O_{i}\right)}{\sum_{j=1}^{|V|}\exp\left(O_{j}\right)},$ (4)
    where $O_{i}^{\prime}$ denotes the output value of the $i$-th node, and $|V|$ represents the count of output nodes, corresponding to the classification categories. Utilizing the Softmax function allows the transformation of output values from multiclassification into a probability distribution, ensuring a sum of 1 within the range [0, 1].
  • Pre-trained Language Model: PLMs undergo initial training using an extensive volume of unlabeled text, enabling them to grasp fundamental language structures such as vocabulary, syntax, semantics, and logic — a phase termed pre-training. Subsequently, this comprehensive language model can be applied to various NLP tasks like machine translation, text summarization, and question-answering systems. To optimize its performance, models need to be trained a second time on a smaller dataset customized for a specific downstream task — a phase known as fine-tuning. This is the “pre-training and fine-tuning” learning paradigm. We can use a visual example to understand the “pre-training and fine-tuning”, as follows: in martial arts novels, a person who wants to become a martial arts master needs to have a solid foundation of internal martial arts, which can be acquired by training on a large variety of techniques. Then, the person can learn a specific skill, such as a sword or a palm strike, and master it quickly and effectively by applying internal martial arts principles. A large number of studies on PLMs have been built on this paradigm, which introduces different architectures (Devlin et al., 2018; Fedus et al., 2022) (e.g., GPT-2 (Radford et al., 2019) and Bert (Devlin et al., 2018)).
  • Large Language Model: LLMs are trained on massive text corpora with tens of billions (or more) of parameters, such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), PaLM (Chowdhery et al., 2022), and LLaMA (Touvron et al., 2023). The goal of LLMs is to enable machines to understand human commands and adherence to human values. Their hallmark lies in the consolidation of two stages: initial pre-training on a vast general-purpose corpus followed by alignment with human values, rather than transitioning to a different domain. LLMs exhibit remarkable adaptability compared to PLMs, transitioning from specialized to general-purpose models. The substantial increase in model size, dataset volume, and computational prowess has resulted in significant enhancements across various tasks and unveiled remarkable capabilities absent in smaller models. For example, GPT-3 has the capability to leverage contextual information, a functionality that GPT-2 (Radford et al., 2019) lacks. This means that when GPT-3 is prompted with task-related examples, it utilizes them to enhance its problem-solving capabilities. The number of parameters for LLMs typically exceeds a hundred billion, and the training data is usually in the range of a few hundred GB to a few TB. A concrete example is that there are multiple versions of the GPT-2 model, with the largest version having 1.5 billion parameters and using 40GB of text data for training. In contrast, the largest version of the GPT-3 model has 175 billion parameters and uses 570GB of text data for training. This example illustrates the significant discrepancy between LLMs and PLMs concerning parameter count and training data volume.
2.2 Factors Propelling Large Language Models
In the past few years, various factors have played a significant role in the swift advancement of LLMs:
  • Data Diversity: data diversity has emerged as a crucial catalyst for the development of LLMs. Recent years have witnessed a surge in the availability and diversity of extensive internet-based data sources, furnishing these models with training data containing rich linguistic and worldly insights. The quality and origins of this data significantly influence the efficacy and capacity of LLMs. Earlier studies trained language models on a single domain of text, such as news articles (Jozefowicz et al., 2016), Wikipedia (Merity et al., 2016), or fiction books (Kiros et al., 2015). However, the proliferation of vast and heterogeneous text corpora across the web—including blogs, social media, forums, reviews, among others—has spurred researchers to pivot towards training models on multi-domain texts. This shift enhances the model’s aptitude for generalization and multitasking. An exemplary instance is Meta AI’s open-source large language model, LLaMA (Touvron et al., 2023), trained on publicly accessible datasets like CommonCrawl [^1] https://commoncrawl.org/get-started, C4 (Raffel et al., 2020), Github [^2] https://github.com/, Wikipedia [^3] https://huggingface.co/datasets/wikipedia, Books (Gao et al., 2020), ArXiv [^4] https://arxiv.org/, and StackExchange [^5] https://stackexchange.com/.
  • Computational Advancement: LLMs are neural network models with a large number of parameters. With the progression of deep learning (DL) technology in recent years, the scale and effectiveness of LLMs have notably escalated. Simultaneously, they present serious challenges due to their substantial computational demands (Kaplan et al., 2020). Computational power signifies a computer system’s capability to execute computational tasks, often quantified in terms of floating-point operations per second (Kaplan et al., 2020). According to OpenAI calculations [^6] https://openai.com/research/ai-and-compute, since 2012, the amount of computation used for global AI training has shown exponential growth, doubling every 3.43 months on average, and now the amount of computation has expanded 300,000 times, far outpacing the rate of computational power growth. Fortunately, hardware innovators like Nvidia continually invent specialized devices such as GPUs, TPUs, and NPUs with amplified computational prowess, LLMs can be trained quickly. For instance, the launch of GPT-3 by OpenAI in May 2020 (Brown et al., 2020) underscores this reliance on robust hardware. Training GPT-3 on a single NVIDIA Tesla V100 GPU would require an estimated 355 years. Conversely, leveraging 1024×A100 GPUs, researchers estimated the potential to train GPT-3 in as little as 34 days. This exemplifies the indispensability of potent hardware in facilitating the development and training of LLMs.
  • Algorithmic Innovation: algorithmic innovation stands as the pivotal driving force behind LLMs, shaping their structure and functionality. From the initial rule- and statistics-based approach, to the later DL-based approach, the algorithms of language models have continuously evolved and improved. Presently, all LLMs are based on the transformer architecture (Vaswani et al., 2017), which employs a self-attention mechanism (see discussion in Section 3). This architecture offers distinct advantages over traditional recurrent (Mikolov et al., 2010) and convolutional neural networks (LeCun et al., 1998), enabling fully parallel computation and adeptly capturing long-distance dependencies. Moreover, there exist numerous iterations and enhancements of the transformer, such as Transformer-XL (Dai et al., 2019), XLNet (Yang et al., 2019), ALBERT (Lan et al., 2019), each targeting specific aspects for optimization. These variations aim to enhance the attention mechanism, refine pre-training objectives, and curtail parameter counts, thereby advancing the transformer architecture in diverse dimensions.
3. PRINCIPLES OF LARGE LANGUAGE MODELS
This section strives to clarify the fundamental principles of LLMs, intending to provide general practitioners with a more profound understanding of the operational mechanisms employed by these models. Fundamental to LLMs is the utilization of deep neural networks to grasp the statistical patterns and semantic nuances of language, enabling the understanding and generation of language. The GPT series, a notable example among these models, incorporates the transformer architecture (Vaswani et al., 2017) and utilizes autoregression to predict consecutive words. The following discussion will concentrate on the GPT-3 (Brown et al., 2020) as an exemplar, providing insights into multiple facets that elucidate the principles of LLMs, as illustrated in Figure 3.
Figure 3: The GPT-3 model architecture. Note that only 2 transformer layers are illustrated for simplicity and illustrative purposes and the actual model consists of 12 transformer layers.
  • Input: The GPT model architecture takes a sequence of symbol representations $(x_{1},\ldots,x_{n})$ as input. This sequence comprises $n$ words, also known as tokens, with GPT-3 explicitly defining the input sequence length as 2048 words.
  • Encoding: The significance of encoding words into vectors is heightened by the fact that the machine learning model operates on numeric inputs. Within the input embedding layer, the initial step involves constructing a comprehensive vocabulary that encompasses all words and assigning each word a unique position within this established vocabulary. For instance, “a” could be designated as 0, “an” as 1, and so forth. In the case of GPT-3, its vocabulary encompasses 50,257 words. Following this, every word can be transformed into a one-hot vector of 50,257 dimensions, where only a single element has the value of one while the rest are zeros. For example, “a” would be represented as [1,0,0,0,…], and “an” as [0,1,0,0,0,…]. This process leads to the encoding of the input sequence into a matrix denoted as $I_{E}\in\mathbb{R}^{2048\times 50257}$.
  • Embedding: The encoding process above transforms every input word into a one-hot vector of 50,257 dimensions. However, this vector exhibits high sparsity, primarily consisting of zeros, leading to significant space inefficiency. To address this drawback, the subsequent step in the input embedding layer employs a embedding matrix, denoted as $W_{E}$, to condense these 50,257-dimensional input vectors into shorter numeric vectors of length $n$ ($n$ is set to 12,288 for GPT-3). This method condenses information about word meanings into a more compact space, effectively reducing the vector’s length. Specifically, the transformation is represented as: $X_{WordEmbedding}=I_{E}\times W_{E}$, where $X_{WordEmbedding}\in\mathbb{R}^{2048\times 12288}$, $I_{E}\in\mathbb{R}^{2048\times 50257}$ and $W_{E}\in\mathbb{R}^{50257\times 12288}$.
  • Positional Encoding: The position and order of elements in a sequence are very important for understanding and generating sequences. Demonstrated through the sentences, “He is a good person and does not do bad things” and “He is a bad person and does not do good things”, a slight alteration in word order - the transition from “good” to “bad” - results in a significant change in the sentence’s semantic meaning. Within the architectural framework of the GPT model, the positional encoding layer serves as a complement to the input embedding layer. This augmentation facilitates the transformer layer in comprehending both the positional and sequential cues within the data. GPT-3 employs the following implementation for positional encodings:
    $PE(\text{pos},\text{ 2i})=\sin\left(\frac{\text{pos}}{10000^{2i/d_{\text{model}}}}\right)$ (5)
    $PE(\text{pos},\text{ 2i+1})=\cos\left(\frac{\text{pos}}{10000^{2i/d_{\text{model}}}}\right)$ (6)
    where $pos$ is the position of the current word within the sentence (e.g. in the sentence “A survey of Large Language Model”, $pos_{A}$=0, $pos_{Large}$=3), $d_{model}$ is the length of the embedding vector 12288 and $i$ is the dimension. However, why delve into such intricate computations for positional encodings? Let’s explore a more straightforward approach to conceptualize it. Initially, considering a text of length $T$, the most basic approach of the positional encoding is counting. This method utilizes $[0,1,2...T-1]$ as positional encodings for each word in the text. For instance, the positional encodings for the $2$-th word would be $[1,1,...,1]$, aligning in length with the embedding vector. However, this encoding strategy presents a notable drawback: within longer sentences, subsequent words hold considerably larger positional encoding values compared to the positional encodings of the initial words. As a result, this disparity inevitably introduces numerical biases in the merged features when combined with word embeddings. To address this issue, normalization becomes a viable consideration. The most straightforward normalization method involves dividing directly by $T$, resulting in positional encodings such as $[0,\frac{1}{T},\ldots,\frac{T-1}{T}]$ (e.g. the 2nd word’s positional encodings=$[\frac{1}{T},\frac{1}{T},\ldots,\frac{1}{T}]$). While this method restricts all positional encodings to the [0,1] range, it introduces a significant inconsistency: the positional values are contingent on the text’s length, causing the positional distinctions between adjacent words in shorter texts to differ from those in longer texts. As illustrated in Figure 4, the positional encodings discrepancy between adjacent word positions in sentence ① is 0.5, while in sentence ②, neighboring word positions are coded with a difference of 0.125. Consequently, neither of these methods proves to be satisfactory. To address this issue, Google devised an alternative approach involving trigonometric functions (Vaswani et al., 2017). As a result, a positional encoding matrix denoted as $X_{PositionalEncoding}\in\mathbb{R}^{2048\times 12288}$.
    Figure 4: Incorrect position encoding.
  • Input Matrix: The input sequence $(x_{1},\ldots,x_{n})$ undergoes processing through both the input embedding layer and the positional embedding layer to generate the input matrix. Specifically, input matrix $X$ $=$ embedding matrix $X_{WordEmbedding}$ $+$ positional encoding matrix $X_{PositionalEncoding}$, where $X\in\mathbb{R}^{2048\times 12288}$, $X_{WordEmbedding}\in\mathbb{R}^{2048\times 12288}$ and $X_{PositionalEncoding}\in\mathbb{R}^{2048\times 12288}$.
  • Masked Multi-Head Self-Attention: Masked multi-head self-attention involves combining a masking technique and a multi-head mechanism within the framework of self-attention. The following discussion commences with self-attention before delving into masked multi-head self-attention for enhanced comprehension.
    Self-Attention: Attention is a crucial cognitive function for human beings, allowing us to concentrate on significant aspects within a complex environment. Within the transformer layer, self-attention mirrors this cognitive functionality by forecasting which input words warrant attention for every output in the sequence, subsequently assessing their significance. For instance, when analyzing the input “I am very” and predicting the output “happy”, the relative contributions of the three words “I am very” to the output “happy” might be quantified as 0.7, 0.1, and 0.2, respectively. The specific steps are as follows:
    Step 1: Initiation commences with the creation of three projection matrices $W_{q}\in\mathbb{R}^{12288\times 128}$, $W_{k}\in\mathbb{R}^{12288\times 128}$, and $W_{v}\in\mathbb{R}^{12288\times 128}$. These matrices are subsequently multiplied with the input matrix $X\in\mathbb{R}^{2048\times 12288}$ to obtain three different matrices $Q$, $K$, and $V$, which respectively represent the query, key, and value.
    Step 2: Following this, a matrix is derived using the formula $Softmax\leavevmode\nobreak\ (Q\times K^{T})$. This matrix then multiplied with $V$ to obtain the output $Y$, represented as $Y=Softmax\leavevmode\nobreak\ (Q\times K^{T})\times V$, where $Q\in\mathbb{R}^{2048\times 128}$, $K^{T}\in\mathbb{R}^{128\times 2048}$, $V\in\mathbb{R}^{2048\times 128}$, and $Y\in\mathbb{R}^{2048\times 128}$. An illustrative analogy can be drawn to searching for a song within a music software - here, $Q$ denotes the queried song’s name, $K$ encompasses all song names within the software database, and $V$ includes the corresponding audio data of all songs in the software. Thus, $Y$ signifies the audio data of the sought-after song.
    Figure 5: Self-Attention and Musk Self-Attention: a normal self-attention block allows a position to attend to all words to its right, while masked self-attention prevents a position from attending to words that come after it in the sequence.
    Masked Self-Attention: Masked self-attention is a modified version of self-attention designed to prevent the model from accessing future words during the prediction of the next word by incorporating a masking technique. For example, with the corpus “abcde”, we can train GPT to predict the next word four times by feeding in “a”, “ab”, “abc” and “abcd” separately. However, this involves segmenting and inputting the sentence four times. The mask technique trick streamlines this process, enabling the model to receive the entire sentence at once and generate an output for each word. The challenge arises from the model peeking at subsequent words, which necessitates the use of masking. When predicting the next word of “a” GPT will only see “a”, and when predicting the next word of “ab”, it will only see “ab”. It’s essential to differentiate between self-attention (utilized in BERT (Devlin et al., 2018)) and masked self-attention (utilized in GPT-3 (Brown et al., 2020)). Figure 5 provides a visual illustration.
    Multi-Head Self-Attention: Multi-head self-attention subdivides the self-attention mechanism into $h$ independent heads to compute in parallel (for GPT-3, $h$ is 96). Each head focuses on different aspects and connections within the sequence, concatenating their outcomes afterward. This approach facilitates the model to collectively concentrate on information derived from diverse representation subspaces, thereby enhancing its capacity for expression. Each head possesses its distinctive projection matrix, each with dimensions $d=d_{model}/h=12288/96=128$. The outcome from each head is symbolized as $Y_{s}$, and when these individual head outputs are combined through concatenation ($96Y_{s}=12288=96\times 128$), the resultant amalgamation is denoted as $Y$.
  • Feedforward Neural Net: In addition to masked multi-head self-attention sub-layers, each of the transformer layers in the GPT model architecture contains a fully connected feedforward neural net. This neural network serves as a mechanism designed to nonlinearly transform its inputs, thereby enhancing the model’s overall expressive capabilities.
  • Add & Norm: The add & norm modules are incorporated into each transformer layer on the GPT model architecture, where “add” represents the skip connection (He et al., 2016) and “norm” represents layer normalization (Ba et al., 2016). The skip connection helps alleviate the issue of gradient vanishing in deep models, enabling the transformer layer to be stacked to a considerable depth. Additionally, layer normalization serves as a regularization strategy to prevent network overfitting. It calculates the mean and variance on each sample independently, without considering other data, thereby mitigating the impact of varying batch sizes.
  • Decoding & Output: After 12 transformer layers, the input matrix $X\in\mathbb{R}^{2048\times 12288}$ transforms into an information-rich matrix $Y\in\mathbb{R}^{2048\times 12288}$, representing the final representation of the input sequence. To achieve the desired result, we employ $Softmax(Y\cdot W)$, where $W\in\mathbb{R}^{12288\times 50257}$, with 50,257 as the vocabulary size. This process generates probabilities assigned to each word in the vocabulary, enabling the selection of the most likely output as GPT’s prediction for each word.
4. APPLICATIONS OF LARGE LANGUAGE MODELS
LLMs have been instrumental across diverse domains, and this section aims to illustrate exemplary applications of LLMs in various fields.
Drug Discovery: The field of drug discovery is characterized by substantial challenges and elevated costs. The utilization of computational methodologies, such as AI, DL, and quantum mechanical techniques, holds the promise of accelerating the identification of potential drug candidates and predicting their properties (Sadybekov & Katritch, 2023; Gorgulla et al., 2022). Among recent advancements in this arena, the integration of LLMs, such as ChatGPT, emerges as a valuable addition, adept at generating and analyzing natural language texts (Savage, 2022). LLMs contribute significantly to the initial phase of drug discovery, particularly in identifying suitable drug targets for diverse diseases (Haley & Roudnicky, 2020; Paananen & Fortino, 2020). They offer preliminary insights into the structure and functionality of protein-based drug targets, subject to further validation by alternative methodologies. Furthermore, LLMs play a pivotal role in subsequent drug discovery steps, aiding in the design and screening of small molecules capable of interacting with these drug targets (Savage, 2022).
Finance: LLMs are driving significant advancements in the finance industry, with significant financial applications including trading and portfolio management (Zhang et al., 2020), financial risk modeling (Mashrur et al., 2020), financial text mining (Gupta et al., 2020; Pagliaro et al., 2021), and financial advisory and customer services (Shah et al., 2020). AI integration in financial advisory and customer services is an emerging and rapidly expanding field. According to (Misischia et al., 2022), AI-driven chatbots already provide more than 37% of supporting functions in various e-service scenarios. Within the financial field, these chatbots are embraced as cost-effective alternatives to human customer service, a trend underscored in the ”Chatbots in Consumer Finance” report [^7] https://www.consumerfinance.gov/data-research/research-reports/chatbots-in-consumer-finance/chatbots-in-consumer-finance. With the advent of LLMs, more and more tasks that were previously considered difficult to handle have become possible, further expanding the potential applications of AI in the financial industry (Li et al., 2023). Notably, there are already some specialized LLMs like BloombergGPT (Wu et al., 2023), a 50-billion-parameter model excelling in financial tasks while maintaining strong performance in general language model benchmarks.
Medical: LLMs are now at the forefront of medical AI with immense potential to improve the efficiency and effectiveness of clinical, educational, and research work (Thirunavukarasu et al., 2023). One example is the performance of LLMs on the MedQA dataset (Jin et al., 2021), which serves as a widely used biomedical question-answering dataset consisting of the United States Medical Licensing Examination (USMLE)-style questions. In this dataset, ChatGPT attains human passing levels and (kung et al., 2023). Additionally, in a span of fewer than six months, Med-PaLM2 achieves proficiency levels close to human experts (Singhal et al., 2023). The strong performance of GPT-4 (OpenAI, 2023) and Med-PaLM2 in medical tests suggests that LLMs may be useful teaching tools for students currently attaining a lower level in such tests (Nori et al., 2023). GPT-4’s meta-prompt feature allows users to explicitly describe the desired role for the chatbot to take on during conversation; a useful example is a “Socratic tutor mode” [^8] https://sites.highlands.edu/tutorial-center/tutor-resources/online-tutor-training/module-4/the-socratic-method/, which encourages students to think for themselves by pitching questions at decreasing levels of difficulty until students can work out solutions to the fuller question at hand. This mode can foster critical thinking and problem-solving skills among students (Thirunavukarasu et al., 2023). ChatGPT demonstrates superior performance in tasks not necessitating specialized knowledge or when such expertise is available in user prompts (Arora et al., 2023). For example, consider discharge summaries as a case in point - they entail repetitive tasks involving information interpretation and compression, often requiring minimal specialized knowledge and user prompts (Touvron et al., 2023).
Legal: LLMs exhibit robust capabilities across various legal tasks, particularly in case prediction and generating legal texts. For instance, ChatGPT achieved a 50.3% accuracy rate on NCBE MBE practice exams, significantly surpassing the 25% random guess level, and achieved passing scores in both Evidence and Torts (Bommarito & Katz, 2022). The integration of ChatGPT holds the potential to reshape the entire legal industry. An Australian law firm experimented with ChatGPT to draft a statement of claim based on the 1992 Mabo case, yielding results akin to those of a first-year lawyer [^9] How ChatGPT and other new AI tools are being used by lawyers, architects and coders - ABC News . Iu and Wong (2022) investigation explored the possibility that ChatGPT could replace litigators by evaluating its drafting and research capabilities. While recommending ChatGPT as a complement to litigators rather than a complete replacement, this suggestion stemmed from ChatGPT’s limitations in meeting the precise accuracy and timeliness demands of legal texts, especially in ensuring continuous access to updated legal research. More recently, advancements such as DeLilegal [^10] https://data.delilegal.com/lawQuestion addressed this issue by integrating with external law databases to enhance retrieval and minimize hallucination. Meanwhile, LLMs in the legal field have emerged in recent months such as ChatLaw (Cui et al., 2022).
5. DRAWBACKS AND FUTURE DIRECTIONS OF LARGE LANGUAGE MODELS
While LLMs offer significant advantages in improving both professional and personal aspects of life, it is essential to acknowledge their accompanying drawbacks. In the subsequent discussion, we will identify and explore some of these limitations inherent in current LLMs, which concurrently provide insights into potential directions for future advancements.
Fairness: LLMs are widely employed for decision-making across various domains, impacting individuals and society. However, these decisions might exhibit bias against marginalized populations. For instance, when utilized in screening resumes for programming positions, LLMs might exhibit a bias favoring male candidates, demonstrating evident gender discrimination. This bias stems from the training of LLMs on extensive and unstructured Internet data, where they inherit stereotypes, misrepresentations, derogatory language, and exclusionary behaviors. These aspects disproportionately affect already disadvantaged and marginalized groups (Bender et al., 2021; Dodge et al., 2021; Sheng et al., 2021). The commonly employed strategy to mitigate bias in LLMs is to remove biased data from the training dataset. However, this strategy does not entirely eliminate bias and often diminishes the effectiveness of language modeling (Meade et al., 2021). Saxena et al. (2020) discuss that the root causes of these problems lie not only in technology but also in socially acceptable definitions of fairness and meaningful interventions to ensure the long-term well-being of all groups. Researchers should consider the needs of disadvantaged groups from the outset and design these technologies proactively, rather than simply reacting to the biases present in their designed systems (Gallegos et al., 2023). The documented bias in LLMs, as highlighted by studies like (Blodgett & O’Connor, 2017Mei et al., 2023Mozafari et al., 2020), underscores the imperative for future efforts in this direction.
Safety: LLMs have diverse applications in industries such as finance and healthcare. However, these applications also present significant security challenges. LLMs may generate outputs that diverge from the provided context, user input, or factual knowledge, a phenomenon referred to as “hallucination” (Bangs et al., 2023; Huang et al., 2023). This undermines their reliability in real-world scenarios, emphasizing the vital necessity for precision, particularly in critical domains like medicine (Dash et al., 2023; Umapathi et al., 2023). As a primary way to avoid these problems, well-aligned LLMs can be developed by including humans in the training loop and using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Glaese et al., 2023). To improve model safety, it is also important to include safety-related cues in the RLHF process, as shown in GPT-4 (OpenAI, 2023). However, RLHF relies heavily on high-quality human feedback data from professional annotators, which is costly and time-consuming. As a result, enhancing the RLHF framework is essential to alleviate the burden on human annotators. Additionally, exploring more efficient annotation methods with assured data quality, such as Large Language Models (LLMs), can assist in the annotation process and enhance overall safety.
Intellectual Property: LLMs possess the capability to create content resembling human-generated material, potentially encroaching upon users’ intellectual property rights. These models rely on training and optimizing using users’ textual data, which may encompass personal information, proprietary knowledge, patented technology, and more. Notably, AI-generated code exemplifies this issue. Code-generation models trained on open-source repositories can produce programs resembling existing ones, potentially disregarding relevant licenses (Yu et al, 2023). For instance, legal action has been taken against Microsoft, GitHub, and OpenAI, alleging copyright infringement due to Copilot reproducing licensed code without adherence to licensing terms [^11] https://www.reuters.com/technology/microsoft-attracting-users-its-code-writing-generative-ai-software-2023-01-25/  [^12] https://www.reuters.com/legal/litigation/openai-microsoft-want-court-toss-lawsuit-accusing-them-abusing-open-source-code-2023-01-27/. Consequently, there’s a pressing need for heightened attention and regulation concerning intellectual property rights concerning LLMs. This is essential to safeguard the legitimate rights and interests of original creators, users, and the public while fostering and respecting the innovation and advancement of LLMs technology (Li et al., 2023).
6. CONCLUSIONS
In recent times, Large Language Models (LLMs) have garnered significant attention from the fields of science, industry, society, and beyond. However, despite the substantial advantages that LLMs offer in enhancing both professional and personal aspects of life, a limited understanding among general practitioners regarding the background and principles of these models impedes their full potential. In light of this, the survey at hand systematically delves into the evolution of LLMs, presenting key concepts and techniques for a comprehensive understanding of these models. Moreover, the survey sheds light on the limitations of LLMs and identifies promising directions for future research. Notably, setting itself apart from other surveys on LLMs, this study prioritizes empowering practitioners, irrespective of their background knowledge, to ensure their proficient utilization in both scientific research and daily tasks.